Equal charges of 3.00x10-9 C are situated at the three corners of a square of side 5.20 m. Find the potential at the unoccupied corner.


1
Expert's answer
2021-04-26T17:11:43-0400

Electric potential created at distance "r" from point charge "q" is "\varphi = \frac{k q}{r}", where "k = 9 \cdot 10^9 \frac{N m^2}{C^2}" is Coulomb's constant. Electric potential created by multiple point charges is just the sum of potentials, created by each charge.

Hence, electric potential at the unoccupied corner is "\varphi = \frac{k q}{a} + \frac{k q}{a} + \frac{k q}{\sqrt{2} a} = \frac{k q}{a}(2+ \frac{1}{\sqrt{2}})", where "a" is the side of a square. Substituting "a = 5.2 m" and "q = 3 \cdot 10^{-9} C", obtain "\varphi \approx 14.06 V".